A friend of yours is considering two cell phone service providers. Provider A charges Rs 120 per month for the service regardless of the number of phone calls made. Provider B does not have a fixed service fee but instead charges Rs 1 per minute for calls. Your friend’s monthly demand for minutes of calling is given by the equation Q(D) = 150 - 50P, where P is the price of a minute. 


a. With each provider, what is the cost to your friend of an extra minute on the phone?

b. In light of your answer to (a), how many minutes would your friend talk on the phone with each provider?

c. How much would he end up paying each provider every month?

d. How much consumer surplus would he obtain with each provider? 

e. Which provider would you recommend that your friend choose? Why? 




1
Expert's answer
2021-03-02T09:45:16-0500

(a) cost of an extra minute in

Provider A=0

Provider B=1

(b) provider A

"Q=150-50p"

"50p=150-Q"

"P=3-Q\/50"

"0=3-Q\/50"

"Q=150"

"150 minutes"

Provider B

"1=3-Q\/50"

"Q\/50=3-1"

"Q=100"

"100minutes"

(C) with provider A

"120 Rs per month"

With provider B

"100*1=100"

"100 Rs per month"

(d) with provider A

"150*3\/2=225-120=105"

With provider B

"100*2\/2=100"

(e) provider A. Because the option offers more consumer surplus.